a certain metal is 30% tin. how many kilograms of this metal must be mixed with 70 kg of a metal that is 70% tin to get a metal that is 50% tin?
Answer by mananth(16946)

--------percent ---------------- quantity
Tin I 70 ---------------- 70 Kgs
Tin II 30 ---------------- x Kgs
Mixture 50 ---------------- 70 + x Kgs

70*70+30 x= 50 ( 70 + x )
4900 + 30x = 3500 + 50 x
30 x + -50 x = 3500 - -4900
-20 x = -1400
/ -20
x = 70
70 Kgs of 30 % Tin II